# Graph Quantum Magic Squares and Free Spectrahedra

### v2

@@ 3.3 Quantum Magic Squares as Free Spectrahedra
In this section we review the connection between quantum magic squares and free spectrahedra, following the outline developed in [DDN20].
@@
A closely related viewpoint is given in [MR4680468], where quantum magic squares are identified with the maximal matrix convex set associated to the (centered) Birkhoff polytope. More precisely, in their notation one has $\mathcal{M}^{(n)}_{s}=(B_{N})_{\max}(s)$ via an explicit description in terms of linear inequalities. Our approach here provides an explicit monic LMI (after an affine translation) tailored to the magic relations, which we will later adapt to incorporate graph-commutation constraints.
In what follows, we give an explicit realization of the associated monic linear pencil that encodes the quantum magic square constraints.
This will allow us to extend the construction to the graph-compatible case introduced above.
@@
These are affine equations in the blocks $X_{kl}$. By solving the affine system, we can express the last row and the last column affinely in terms of the $(n-1)^{2}$ independent blocks $\{X_{ij}\}_{1\leq i,j\leq n-1}$:
@@
The inequalities $X_{kl}\succeq 0$ can be grouped into a block-diagonal matrix, yielding the linear pencil
@@
Substituting into (5) gives
@@
With these values, one verifies that $\alpha_{kl}+\frac{1}{n}\sum_{i,j}c_{kl}^{\,ij}=\frac{1}{n}$ for all $(k,l)$.
In summary, the set of quantum magic squares admits a concrete realization as a free spectrahedron, defined by the monic linear pencil $\widehat{L}(Y)$. This description highlights the geometric nature of $\mathcal{M}^{(n)}$ as a compact matrix convex set, whose structure is completely determined by finitely many linear matrix inequalities. In the next section, we extend this framework to the graph-setting by introducing graph quantum magic squares, obtained by adding commutation constraints with the adjacency matrix of a graph $\Gamma$.
@@ 3.4 Graph Quantum Magic Squares as Free Spectrahedra
We now extend the previous construction to the case of graph quantum magic squares. The key idea is to include the commutation relations with the adjacency matrix of the graph into the affine system and the resulting linear pencil.
Our main structural result is that for $k$-regular graphs, this set admits an explicit (monic) linear matrix inequality representation.
@@
In the following subsections, we construct this pencil explicitly, first for the simple case of complete graphs, and then for general $k$-regular graphs.
@@ 3.4.1 Complete graphs
is automatically satisfied. Indeed, the magic relations imply that both row and column sums are the identity, hence
@@
Therefore, for $K_{n}$ the commutation condition adds no further restriction:
@@
and the free-spectrahedral LMI coincides with the one obtained above.
@@ 3.4.2 Affine parametrization under commutation for $k$-regular graphs
Let $\Gamma=\bigcup_{t=1}^{N}\Gamma_{t}$ be a $k$-regular graph on $n$ vertices with adjacency matrix $A_{\Gamma}$. Let $B\in\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))$ be a block matrix commuting with $A_{\Gamma}\otimes I_{s}$. Then there exist Hermitian matrices $\Lambda_{R}^{(t)},\Lambda_{C}^{(t)}\in\operatorname{Mat}_{n}(\operatorname{Her}_{s}(\mathbb{C}))$ such that, for every $t=1,\dots,N$,
$\sum_{j=1}^{n}B_{ij}=\Lambda_{R}^{(t)}\quad(\forall i\in\Gamma_{t})\qquad\text{and}\qquad\sum_{i=1}^{n}B_{ij}=\Lambda_{C}^{(t)}\quad(\forall j\in\Gamma_{t}).$
@@
Let $\mathbf{1}_{t}\in\mathbb{R}^{n}$ be the indicator vector of the vertex set of the component $\Gamma_{t}$. Since each $\Gamma_{t}$ is connected and $k$-regular, $A_{\Gamma}\mathbf{1}_{t}=k\,\mathbf{1}_{t}$. The condition $B(A_{\Gamma}\otimes I_{s})=(A_{\Gamma}\otimes I_{s})B$ implies that for any vector $v\in\mathbb{C}^{s}$,
$k\,B(\mathbf{1}_{t}\otimes v)=B(A_{\Gamma}\otimes I_{s})(\mathbf{1}_{t}\otimes v)=(A_{\Gamma}\otimes I_{s})B(\mathbf{1}_{t}\otimes v).$
Thus, $B(\mathbf{1}_{t}\otimes v)$ must lie in the $k$-eigenspace of $A_{\Gamma}\otimes I_{s}$, which is spanned by $\{\mathbf{1}_{r}\otimes w\mid r=1\dots N,w\in\mathbb{C}^{s}\}$. Since the graph components are disjoint, one can verify that the action remains localized on the component, implying $B(\mathbf{1}_{t}\otimes v)=\mathbf{1}_{t}\otimes(T_{t}v)$ for some linear map $T_{t}$. In block form, this means $\sum_{j=1}^{n}B_{ij}=T_{t}=:\Lambda_{R}^{(t)}$ for all $i\in\Gamma_{t}$. An analogous argument yields the column sums. ∎
Ideally, the magic relations require all row and column sums to be equal to $I_{s}$. However, thanks to Proposition 3.10, for a matrix in the commutant of $A_{\Gamma}$, it is sufficient to impose this condition once per connected component. Specifically, setting $\Lambda_{R}^{(t)}=I_{s}$ (and similarly for columns) fixes the sum for all rows in that component. Consequently, imposing the magic relations removes exactly $N$ degrees of freedom (where $N$ is the number of connected components) from the $d_{\Gamma}$ independent parameters provided by imposing the commutation constraints.
We can now proceed with the explicit parametrization. Let
@@
be an index set of independent positions for the commutant. In practice, we choose the variables $X_{i^{\prime}j^{\prime}}$ with $(i^{\prime},j^{\prime})\in\mathcal{I}$ from the first and second rows until $d_{\Gamma}$ independent parameters are collected.
@@
The magic constraints allow us to express $N$ of these variables in terms of the others. Let $\mathcal{K}\subset\mathcal{I}$ be the set of $N$ dependent indices. The set of remaining independent indices is $\mathcal{I}_{\mathrm{ind}}=\mathcal{I}\setminus\mathcal{K}$.
@@
Finally, we observe that the constant matrix $X^{\circ}_{ij}=\tfrac{1}{n}I_{s}$ satisfies both commutation and magic constraints. Indeed, for any $k$-regular graph, the all-ones matrix $J$ commutes with $A_{\Gamma}$ (since $A_{\Gamma}J=JA_{\Gamma}=kJ$), and thus the scalar matrix $\tfrac{1}{n}I_{s}$ is a valid solution. This implies that the constant term in our parametrized pencil corresponds to this point, and after rescaling, we obtain the monic pencil stated in Theorem 3.9:
@@ 3.4.3 Composition of affine maps and explicit coefficients
In the graph case, we restrict the variables $X_{ij}$ to the subspace of solutions to the commutation constraints. Using the linear map $\Psi_{\mathrm{com}}$ defined above, we can express every block $X_{ij}$ in terms of the independent variables indexed by $\mathcal{I}$:
@@
Composing this with the affine parametrization for QMS yields
@@
Hence, defining the combined coefficients as
$d_{kl}^{(i^{\prime}j^{\prime})}=\sum_{i,j=1}^{n-1}c_{kl}^{ij}\,r^{\,ij}_{i^{\prime}j^{\prime}},$
we see that the explicit form of the matrices $B_{i^{\prime}j^{\prime}}$ appearing in the LMI of Theorem 3.9 is given by diagonal matrices whose entries are precisely these coefficients $d_{kl}^{(i^{\prime}j^{\prime})}$.
After the normalisation shift
@@
we recover exactly the monic affine parametrization
@@ 3.5 GQMS and Arveson extreme points
Having established in Theorem 3.9 that $\mathcal{M}^{(\Gamma)}$ is affinely isomorphic to a compact free spectrahedron, we may now apply Theorem 2.12 to conclude that
@@
In the non-graph case, [DDN20] showed that every quantum permutation matrix is an Arveson extreme point of the free spectrahedron $\mathcal{M}^{(n)}$. Since graph quantum permutation matrices form a subclass of quantum permutation matrices subject only to additional linear constraints, the same argument used in [DDN20] also applies in the graph setting.
This yields the following corollary.
@@ 3.5.1 Concluding Remarks on the Free-Spectrahedral Setting
The set $\mathcal{M}^{(\Gamma)}$ is therefore affinely isomorphic to a compact free spectrahedron, defined by finitely many linear matrix inequalities with scalar coefficients. By the general theory of free spectrahedra (see Theorem 2.12), its matrix convex hull is generated by its Arveson extreme points.
Our corollary 3.11 shows that every graph quantum permutation matrix belongs to the set $\operatorname{ArvesonExt}(\mathcal{M}^{(\Gamma)})$. Consequently, in analogy with the non-graph case [DDN20]*Corollary 20, the extremal structure of $\mathcal{M}^{(\Gamma)}$ is strictly richer than that of $\mathcal{P}^{(\Gamma)}$ whenever the Birkhoff–von Neumann property fails for $\Gamma$.
Observation 3.12 ($\mathcal{M}^{(n)}$ and $\mathcal{M}^{(\Gamma)}$ over a unital $C^{*}$-algebra).
If we take any unital $C^{*}$-algebra $\mathcal{A}$, all previous results remain valid upon replacing $\operatorname{Her}_{s}(\mathbb{C})$ with $\mathcal{A}_{\mathrm{sa}}$ and $I_{s}$ with $I_{\mathcal{A}}$.
@@ 4 Future Directions
The results obtained in this work suggest several directions for further investigation.
4.1 (1) Beyond $C_{4}$: the (Quantum) Birkhoff–von Neumann problem for graphs
Our construction and separation argument for $C_{4}$ raise the broader question of whether the Birkhoff–von Neumann theorem fails for other graphs, in the quantum setting. A natural class of candidates includes vertex-transitive or symmetric graphs, such as the cycle $C_{5}$, the Petersen graph, and more generally any $k$-regular graph with large automorphism group.
This leads to the conjectural extension:

### v3

@@ 3.3 Quantum Magic Squares as Free Spectrahedra
In this section, we study quantum magic squares as free spectrahedra, following [DDN20].
@@
A closely related point of view is given by [BNS23], where quantum magic squares are identified with the maximal matrix convex set associated to the (centered) Birkhoff polytope. More precisely, in their notation $\mathcal{M}^{(n)}_{s}=(B_{N})_{\max}(s)$ via an explicit description in terms of linear inequalities. Our approach here provides an explicit (after an affine translation) monic linear pencil adapted to magic relations, which we will later use to enforce commutation with the adjacency matrix of a graph.
In what follows, we give an explicit monic linear pencil for the quantum magic square.
@@
These are affine equations in the blocks $X_{kl}$. By solving the affine system, we can express the last row and the last column in terms of the $(n-1)^{2}$ independent blocks $\{X_{ij}\}_{1\leq i,j\leq n-1}$:
@@
The inequalities $X_{kl}\succeq 0$ can be put into a block-diagonal matrix, yielding the linear pencil
@@
and substituting it into (6) gives us
@@
With these values, we can verify that $\alpha_{kl}+\frac{1}{n}\sum_{i,j}c_{kl}^{\,ij}=\frac{1}{n}$ for all $(k,l)$.
In summary, the set of quantum magic squares admits a concrete description as a free spectrahedron, defined by the monic linear pencil $\widehat{L}(Y)$. This description highlights the geometry of $\mathcal{M}^{(n)}$ as a compact matrix convex set: it is completely determined by finitely many linear matrix inequalities. In the next section, we extend this framework to the set of graph quantum magic squares.
@@ 3.4 Graph Quantum Magic Squares as Free Spectrahedra
We now extend the previous construction to the case of graph quantum magic squares. The key idea is to encode the commutation relations with the adjacency matrix of the graph into the affine system and the resulting linear pencil.
For $k$-regular graphs, this set admits an explicit (monic) linear matrix inequality representation.
@@
In the next subsections, we construct this pencil explicitly, first for the simple case of complete graphs, and then for general $k$-regular graphs.
@@ 3.4.1 Complete graphs
is automatically satisfied. Indeed, the magic relations imply that both row and column sum to the identity, hence
@@
Therefore, for $K_{n}$ the commutation condition does not add any further restriction. We, hence, have
@@
and the LMI coincides with the one obtained above.
@@ 3.4.2 Affine parametrization under commutation for $k$-regular graphs
Let $\Gamma$ be a $k$-regular graph with $N$ connected components $\Gamma=\bigcup_{t=1}^{N}\Gamma_{t}$ and let $A_{\Gamma}$ be its adjacency matrix.
Let $X=\left(X_{ij}\right)_{i,j=1}^{n}\in\mathrm{Mat}_{n}(\mathrm{Her}(s))$ be such that
$X(A_{\Gamma}\otimes I_{s})=(A_{\Gamma}\otimes I_{s})X.$
Then there exist Hermitian matrices $\Lambda_{R}^{(t)},\Lambda_{C}^{(t)}\in\mathrm{Her}(s)$ such that, for every $t=1,\dots,N$,
$\sum_{j=1}^{n}X_{ij}=\Lambda_{R}^{(t)}\quad(\forall i\in V_{t}),\qquad\sum_{i=1}^{n}X_{ij}=\Lambda_{C}^{(t)}\quad(\forall j\in V_{t}),$
where $V_{t}$ is the set of vertices of the component $\Gamma_{t}$.
Before proceeding with the proof we recall a property of the eigenspaces of a $k$-regular graph.
Observation 3.11.
If $\Gamma$ is $k$-regular with $n$ vertices, then $A_{\Gamma}\mathbf{1}=k\mathbf{1}$, where $\mathbf{1}=(1,\dots,1)^{T}\in\mathbb{C}^{n}$. In particular, if $\Gamma=\bigsqcup_{t=1}^{N}\Gamma_{t}$ is a $k$-regular graph with $N$ connected components, it also holds that
$A_{\Gamma}\mathbf{1}_{t}=k\,\mathbf{1}_{t}\qquad\forall t=1,\dots,N,$
where $\mathbf{1}_{t}$ is the $n$-dimensional vector defined by
$(\mathbf{1}_{t})_{i}=\begin{cases}1,&i\in V(\Gamma_{t}),\\ 0,&i\notin V(\Gamma_{t}),\end{cases}\qquad i=1,\dots,n.$
Furthermore, since each $\Gamma_{t}$ is connected and $k$-regular, the eigenspace relative to the eigenvalue $k$ of $A_{\Gamma}$ is $\ker(A_{\Gamma}-kI)=\mathrm{span}\{\mathbf{1}_{t}\mid t=1,\dots,N\}$. So the eigenspace related to the eigenvalue $k$ of $A_{\Gamma}\otimes I_{s}$ is
$\ker(A_{\Gamma}\otimes I_{s}-kI)=\mathrm{span}\{\mathbf{1}_{t}\otimes e_{p}\mid t=1,\dots,N,\;p=1,\dots,s\},$
where $\{e_{1},\dots,e_{s}\}$ is the canonical basis of $\mathbb{C}^{s}$.
@@
Let $V=\{1,\dots,n\}$ be the set of vertices of $\Gamma=\bigsqcup_{t=1}^{N}\Gamma_{t}$, and $V_{t}$ the set of vertices of $\Gamma_{t}$.
Now fix an index $t_{0}\in\{1,\dots,N\}$ and a vector $v\in\mathbb{C}^{s}$. From the commutation $X(A_{\Gamma}\otimes I_{s})=(A_{\Gamma}\otimes I_{s})X$ it follows that
$(A_{\Gamma}\otimes I_{s})\left[X(\mathbf{1}_{t_{0}}\otimes v)\right]=X(A_{\Gamma}\otimes I_{s})(\mathbf{1}_{t_{0}}\otimes v)=X(k\,\mathbf{1}_{t_{0}}\otimes v)=k\left[X(\mathbf{1}_{t_{0}}\otimes v)\right],$
so $X(\mathbf{1}_{t_{0}}\otimes v)$ is an eigenvector of $A_{\Gamma}\otimes I_{s}$ with eigenvalue $k$. Therefore there exist coefficients $c_{tp}(v)\in\mathbb{C}$ such that
$X(\mathbf{1}_{t_{0}}\otimes v)=\sum_{t=1}^{N}\sum_{p=1}^{s}c_{tp}(v)(\mathbf{1}_{t}\otimes e_{p}).$
The $i$-th component of $X(\mathbf{1}_{t_{0}}\otimes v)$ is
$\left[X(\mathbf{1}_{t_{0}}\otimes v)\right]_{i}=\sum_{j=1}^{n}X_{ij}(\mathbf{1}_{t_{0}})_{j}\,v=\sum_{j\in V_{t_{0}}}X_{ij}v.$ (7)
On the other hand, if $i\in V_{\ell}$ then $(\mathbf{1}_{\ell})_{i}=1$ and $(\mathbf{1}_{t})_{i}=0$ for $t\neq\ell$, so from the previous decomposition we obtain
$\left[X(\mathbf{1}_{t_{0}}\otimes v)\right]_{i}=\sum_{p=1}^{s}c_{\ell p}(v)e_{p}\qquad(\forall i\in V_{\ell}).$ (8)
So, comparing (7) and (8), for each $\ell$ and for each $i\in V_{\ell}$ we have
$\sum_{j\in V_{t_{0}}}X_{ij}v=\sum_{p=1}^{s}c_{\ell p}(v)e_{p}.$
Since the right-hand side depends only on the component index $\ell$ (and the vector $v$), but not on the specific vertex $i$, it follows that the matrix sum $\sum_{j\in V_{t_{0}}}X_{ij}$ is the same for all $i\in V_{\ell}$.
In particular, by summing over all $t_{0}=1,\dots,N$, we obtain that for every $i\in V_{\ell}$:
$\sum_{j=1}^{n}X_{ij}=\sum_{t_{0}=1}^{N}\left(\sum_{j\in V_{t_{0}}}X_{ij}\right)=:\Lambda^{(\ell)}_{R}$,
where $\Lambda^{(\ell)}_{R}\in\mathrm{Mat}_{s}(\mathbb{C})$ depends only on the connected component $V_{\ell}$. Since $X_{ij}\in\mathrm{Her}(s)$, the sum $\Lambda^{(\ell)}_{R}$ is also Hermitian. For the columns, we follow the same reasoning with $X^{T}$ instead of $X$: since $A_{\Gamma}^{T}=A_{\Gamma}$, the commutation also holds,
$X^{T}(A_{\Gamma}\otimes I_{s})=(A_{\Gamma}\otimes I_{s})X^{T}.$
Applying the above to $X^{T}$ we obtain that
$\sum_{i=1}^{n}X_{ij}=:\Lambda^{(\ell)}_{C}\in\mathrm{Mat}_{s}(\mathbb{C}),$
the column sums of $X$ are constant on each connected component $l$, yielding the same matrix $\Lambda^{(\ell)}_{C}$. ∎
In general, magic relations require that all row and column sums be equal to $I_{s}$. However, thanks to Proposition 3.10, for a matrix in the commutant of $A_{\Gamma}$, it is sufficient to impose this condition once per connected component. In particular, imposing $\Lambda_{R}^{(t)}=I_{s}$ (and similarly for the columns) fixes the sum for all rows of that component. Consequently, imposing magic relations removes exactly $N$ degrees of freedom (where $N$ is the number of connected components) from the $d_{\Gamma}$ independent parameters provided by imposing the commutation constraints.
Now we perform the explicit parameterization. Let
@@
be a set of independent position indices for the commutant. In practice, we choose the variables $X_{i^{\prime}j^{\prime}}$ with $(i^{\prime},j^{\prime})\in\mathcal{I}$ from the first and second rows until exactly $d_{\Gamma}$ independent parameters are collected.
@@
Magic constraints allow us to express $N$ of these variables as functions of the others. Let $\mathcal{K}\subset\mathcal{I}$ be the set of $N$ dependent indices and $\mathcal{I}_{\mathrm{ind}}=\mathcal{I}\setminus\mathcal{K}$ be the set of the remaining independent indices.
@@
Finally, we observe that the constant matrix $\left(X^{\circ}_{ij}\right)_{ij}=\left(\tfrac{1}{n}I_{s}\right)_{ij}$ satisfies both commutation and magic constraints. Indeed, for any $k$-regular graph, the all-ones matrix $J$ commutes with $A_{\Gamma}$ (since $A_{\Gamma}J=JA_{\Gamma}=kJ$), and thus the scalar matrix $\tfrac{1}{n}I_{s}$ is a valid solution. This implies that the constant term in our parametrized pencil corresponds to this point, and after rescaling, we obtain the monic pencil of Theorem 3.9:
@@ 3.4.3 Composition of affine maps and explicit coefficients
In the GQMS case, we restrict the variables $X_{ij}$ to the subspace of solutions to the commutation constraints. Using the linear map $\Psi_{\mathrm{com}}$ defined above, we can express every block $X_{ij}$ in terms of the independent variables indexed by $\mathcal{I}$:
@@
After composing this with the affine parametrization for QMS, we get
@@
Hence, defining
$d_{kl}^{(i^{\prime}j^{\prime})}:=\sum_{i,j=1}^{n-1}c_{kl}^{ij}\,r^{\,ij}_{i^{\prime}j^{\prime}},$
we get that the explicit form of $B_{i^{\prime}j^{\prime}}$ in the LMI of Theorem 3.9 is given by diagonal matrices whose entries are precisely the coefficients $d_{kl}^{(i^{\prime}j^{\prime})}$.
Now, if we perform the shift
@@
we get exactly the monic affine parametrization
@@ 3.5 GQMS and Arveson extreme points
Theorem 3.9 establishes that $\mathcal{M}^{(\Gamma)}$ is affinely isomorphic to a compact free spectrahedron. We may now apply Theorem 2.12 to conclude that
@@
In the non-graph case, [DDN20] showed that every quantum permutation matrix is an Arveson extreme point of the free spectrahedron $\mathcal{M}^{(n)}$. Since graph quantum permutation matrices form a subclass of quantum permutation matrices subject only to additional linear constraints, the same argument used in [DDN20] also applies in the graph setting and we have the following corollary.
@@ 3.5.1 Concluding Remarks on the Free-Spectrahedral Setting
The set $\mathcal{M}^{(\Gamma)}$ is affinely isomorphic to a compact free spectrahedron, defined by finitely many linear matrix inequalities with scalar coefficients. By Theorem 2.12, its matrix convex hull is generated by its Arveson extreme points.
Corollary 3.12 shows that every graph quantum permutation matrix belongs to the set $\operatorname{ArvesonExt}(\mathcal{M}^{(\Gamma)})$. Hence, in analogy with the non-graph case [DDN20]*Corollary 20, the set $\mathcal{M}^{(\Gamma)}$ is strictly larger than that the set $\mathcal{P}^{(\Gamma)}$ whenever the Birkhoff–von Neumann property fails for $\Gamma$.
Observation 3.13 ($\mathcal{M}^{(n)}$ and $\mathcal{M}^{(\Gamma)}$ over a unital $C^{*}$-algebra).
If we take any unital $C^{*}$-algebra $\mathcal{A}$, all previous results hold upon replacing $\operatorname{Her}_{s}(\mathbb{C})$ with $\mathcal{A}_{\mathrm{sa}}$ and $I_{s}$ with $I_{\mathcal{A}}$.
@@ 4 Future Directions
These results suggest some avenues for future work.
4.1 Beyond $C_{4}$: the (Quantum) Birkhoff–von Neumann problem for graphs
Having found an example of the failure of the Birkhoff–von Neumann theorem for the graph $C_{4}$, suggests investigating whether the Birkhoff–von Neumann theorem fails in the quantum setting for other graphs. A natural class of candidates is vertex-transitive or symmetric graphs, such as the cycle $C_{5}$, the Petersen graph, and, more generally, any $k$-regular graph with a large automorphism group.
This leads to the conjecture: